# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ 2.1 Isotropic case
Figure 1: Estimation performance of MOSEG (Cho and Owens,, 2024), CHARCOAL (Gao and Wang,, 2022; only applicable when $n>p$), and the proposed methods: McScan (Section 3.1; a variant of Cho et al.,, 2025), QcScan (Section 3.2), OcScan (Section 4.1) and OcScan.R (Section 4.2), in the isotropic setting for varying $\mathfrak{s}$ ($x$-axis) and sample size $n$ (left to right). Results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines correspond to where $\mathfrak{s}=\sqrt{p}$. The $x$-axis is displayed on a log scale, and the $y$-axis on a squared root scale. This is excerpted from Figure 3 in Section 5.
@@ 2.2 General covariance
$\mathsf{E}(Y_{t}|\mathbf{x}_{t})=\left[(\bm{\Sigma}^{1/2})^{\dagger}\mathbf{x}_{t}\right]^{\top}\bm{\Sigma}^{1/2}\left(\bm{\beta}_{0}\cdot\mathbb{I}_{\{t\leq\theta\}}+\bm{\beta}_{1}\cdot\mathbb{I}_{\{t>\theta\}}\right),$ (2)
@@
where $(\bm{\Sigma}^{1/2})^{\dagger}$ denotes the (Moore–Penrose) pseudoinverse of $\bm{\Sigma}^{1/2}$. Hence the lower bounds reported in Table 1 are straightforward generalisation of Theorem 1 to the model class $\mathcal{P}^{\mathfrak{s},n,p}_{\bm{\Sigma},\sigma}(\tau)$ with general $\bm{\Sigma}$. It is readily seen that the phase transition is determined by the sparsity of the covariance weighted differential parameter, i.e. the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. Also, the lower bounds are formulated with the size of the change measured by $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}$, complementing the following lemma:
Lemma 2 (Lemma 1 of Cho et al.,, 2025).
@@
Let $\mathsf{P}_{0}(\bm{\beta})$ denote the joint distribution of $\{(\mathbf{x}_{t},Y_{t}):\,t=1,\ldots,n\}$ such that $Y_{t}=\mathbf{x}_{t}^{\top}\bm{\beta}+\varepsilon_{t}$ for all $t$. We also denote by $\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})$ the joint distribution of $\{(\mathbf{x}_{t},Y_{t}):\,t=1,\ldots,n\}$ under (1). Then the total variation distance between $\mathsf{P}_{0}(\bm{\beta})$ and $\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})$, denoted by $\mathrm{TV}\bigl{(}\mathsf{P}_{0}(\bm{\beta}),\,\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})\bigr{)}$, satisfies
@@
$\frac{1}{100}\leq\frac{\min_{\bm{\beta},\bm{\beta}_{0},\bm{\beta}_{1}:\,\bm{\beta}_{1}-\bm{\beta}_{0}=\bm{\delta}}\mathrm{TV}\bigl{(}\mathsf{P}_{0}(\bm{\beta}),\,\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})\bigr{)}}{\min\left\{1,\,\sqrt{\frac{\theta(n-\theta)}{n\sigma^{2}}\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}}\right\}}\leq\frac{3\sqrt{3}}{2}.$
@@
Figure 2 numerically confirms these observations when $\mathsf{Cov}(\mathbf{x}_{t})=[0.6^{|i-j|}]_{i,j}$. Specifically, the best performance of various change point estimation methods exhibits a steeper dependence on the intrinsic sparsity (middle panel) than on the standard sparsity (left panel), which is in alignment with the theoretical bounds shown in Table 1 (see also the right panel). Besides, we notice that McScan, a method attaining minimax rate-optimality in the sparse regime, has its estimation error exceed that of QcScan, a method tailored to achieve optimal performance in the dense regime, close to where the inherent sparsity $\mathfrak{s}$ meets $\sqrt{p}$ (middle panel). On the other hand, this phenomenon occurs far before the sparsity of $\bm{\delta}$ reaches $\mathfrak{s}_{\delta}=\sqrt{p}$, which is accounted for by that $|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=p$ in all the scenarios considered in the left panel.
@@
Figure 2: Estimation performance of MOSEG (Cho and Owens,, 2024), CHARCOAL (Gao and Wang,, 2022), and the proposed methods: McScan (Section 3.1; a variant of Cho et al.,, 2025), QcScan (Section 3.2), OcScan (Section 4.1) and OcScan.R (Section 4.2), when $\bm{\Sigma}=[0.6^{|i-j|}]_{i,j}$. Taken from (M2) in Section 5, we set $n=300$, $p=200$, $\bm{\beta}_{0}+\bm{\beta}_{1}=\mathbf{0}$ and $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}=2$. In the left, the $x$-axis denotes the standard sparsity $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$ while in the middle, it is the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. In each scenario, results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The right panel plots the quantity $\min(|\bm{\Sigma}^{1/2}\bm{\delta}|_{0},\sqrt{p})$ as either $\mathfrak{s}_{\delta}$ or $\mathfrak{s}$ varies. This quantity is the influencing factor, save logarithmic factors, appearing in the minimax rate of estimation (see Table 1), and represents a proxy of the difficulty of the estimation problem. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$ or $\mathfrak{s}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@
In the high-dimensional change point detection literature, the phase transition with respect to sparsity has been noted in the context of mean change detection, see Enikeeva and Harchaoui, (2019), Liu et al., (2021) and Pilliat et al., (2023), all considering the case of cross-sectional independence. Theorem 1 is a first of its kind for the linear model setting where the cross-sectional dependence is present in $\mathbf{x}_{t}Y_{t}$ (whose expectation is piecewise constant as $\mathsf{E}(\mathbf{x}_{t}Y_{t})=\bm{\Sigma}\bm{\beta}_{0}+\bm{\Sigma}\bm{\delta}\cdot\mathbb{I}_{\{t\geq\theta+1\}}$), even if $\mathsf{Cov}(\mathbf{x}_{t})=\mathbf{I}$. Moreover, the sparsity inherent to this change point problem is characterised by $|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$ rather than the $\ell_{0}$-norm of $\bm{\delta}$, which also sets our findings apart from the existing work.
@@ 3.1 Sparse regime
Cho et al., (2025) propose McScan, a data segmentation method for the multiple change point problem in linear models. Motivated by the observation that $\mathsf{E}(\mathbf{x}_{t}Y_{t})$ contains the information about local parameter vector, it searches for change points by scanning the local covariance between $Y_{t}$ and $\mathbf{x}_{t}$. Avoiding the costly computation of local parameter estimates via $\ell_{1}$-regularision, it is highly efficient computationally and moreover, attains better statistical efficiency over a wider class of change point problems when compared to the state-of-the-art, by relaxing the conditions of exact sparsity imposed on the regression or differential parameters.
@@
We propose to evaluate $\bar{T}_{k}$ at strategically selected locations for the detection and estimation of the change point (if any). For this, we adopt the advanced optimistic search proposed by Kovács et al., (2024), see Algorithm 1 for the generic pseudo code applicable with any detector statistic denoted by $V_{k}$. It first evaluates $V_{k}$ on a dyadic grid denoted by $\mathcal{D}$, for a given trimming parameter $\varpi$, and identifies the maximiser $k^{*}$. If $V_{k}$ does not exceed some threshold $\zeta$, we quit the algorithm while if $V_{k}>\zeta$, we perform a divide-and-conquer-type search through calling the function OS over the interval determined with $k^{*}$ and its neighbours on the dyadic grid. Through this, we iteratively compare $V_{k}$ at dynamically selected search locations based on prior evaluations and discard the sections of the data that are not likely to contain any change point. Replacing the conventional exhaustive search, this adaptive strategy significantly reduces the total computational complexity from linear to logarithmic, while maintaining the statistical efficiency. We apply Algorithm 1 with McScan-specific detector statistic $\bar{T}_{k}$, threshold $\zeta_{\text{\tiny M}}$ and trimming parameter $\varpi_{\text{\tiny M}}$, to obtain the estimators of the number and location of the change point, $(\widehat{\theta}_{\text{\tiny M}},\widehat{q}_{\text{\tiny M}})$. With a slight abuse of terminology, we refer to the resulting method also as McScan (maximum covariance scanning).
@@
Output: $(\widehat{\theta},\,\widehat{q})$
Algorithm 1 Advanced optimistic search (adapted from Kovács et al.,, 2024).
@@
$|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}\Delta\geq\bar{c}_{0}\|\bm{\Sigma}\|\Psi^{2}\log(p\log(n)),$ (5)
or that there is no change point, with $q=0$ and $\theta=n$. Then, McScan applied with $\zeta_{\text{\tiny M}}=\bar{c}\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\log(p\log(n))}$ and $\varpi_{\text{\tiny M}}$, outputs $(\widehat{\theta}_{\text{\tiny M}},\widehat{q}_{\text{\tiny M}})$ which satisfy
$\mathsf{P}\left\{\widehat{q}_{\text{\tiny M}}=q\,;\,|\widehat{\theta}_{\text{\tiny M}}-\theta|\leq\bar{c}_{1}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}\|\bm{\Sigma}\|\Psi^{2}\log(p\log(n))\right\}\geq 1-\bar{c}_{2}(p\log(n))^{-1}$
[…]
@@
Thus, as long as $\sigma_{*}$ is bounded away from zero, McScan is minimax rate optimal up to a logarithmic factor in the sparse regime $\mathfrak{s}<\sqrt{p}$ (cf. Table 1), without requiring any prior knowledge of $\bm{\Sigma}$, or ruling out strong collinearity among the variables that results in rank deficient $\bm{\Sigma}$ – a situation that often arises with increasing number of variables. This is in stark contrast to the most existing methods that require $\mathsf{Cov}(\mathbf{x}_{t})$ to have bounded spectrum or characterise their properties in terms of the smallest eigenvalue of $\bm{\Sigma}$; see Section S1.1 of Cho et al., (2025) and also Figure 5 in Section 5 for numerical validation. When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ as in (2) allows for re-formulating the condition in (5), to
@@
Since $|\bm{\Sigma}^{1/2}\bm{\delta}|_{\infty}\geq\mathfrak{s}^{-1/2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}$ (recall that $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$), the above displayed condition is independent of $\sigma_{*}$, and matches the minimax lower bound on the detection boundary in the sparse regime, provided that $\log(n)\lesssim p$; similar observations are made about the rate of estimation reported in Theorem 3.
@@
The original McScan algorithm of Cho et al., (2025) combines the detector statistic in (3) with the narrowest-over-threshold principle (Baranowski et al.,, 2019) and the seeded intervals (Kovács et al.,, 2023), evaluating the detector statistics over strategically selected intervals and selecting the significant local maximisers for the detection of multiple change points. Theorem 3 is a variant of the more general Theorem 2 in Cho et al., (2025) which, permitting serial dependence and sub-Weibull tail behaviour in the data, shows the consistency of McScan in multiple change point detection, all without placing any condition on the exact sparsity of regression or differential parameters, or that on $p$ beyond what is implied by the detection lower bound comparable to (5). There, it is further demonstrated that McScan allows for a strictly wider class of change point models in comparison with the state-of-the-art.

### v3

@@ 2.1 Isotropic case
Figure 1: Estimation performance of MOSEG (Cho and Owens, 2024), CHARCOAL (Gao and Wang, 2022; only applicable when $n>p$), and the proposed methods: McScan (Section 3.1; a variant of Cho et al., 2025), QcScan (Section 3.2), OcScan (Section 4.1) and OcScan.R (Section 4.2), in the isotropic setting ($\bm{\Sigma}=\mathbf{I}$) for varying $\mathfrak{s}$ ($x$-axis) and sample size $n$ (left to right) while $p=900$ and $|\bm{\delta}|_{2}=8$. Results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines correspond to where $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$. The $x$-axis is displayed on a log scale, and the $y$-axis on a squared root scale. This is excerpted from Figure 3 in Section 5.
@@ 2.2 General covariance
$\mathsf{E}(Y_{t}\mid\mathbf{x}_{t})=\left[(\bm{\Sigma}^{1/2})^{\dagger}\mathbf{x}_{t}\right]^{\top}\bm{\Sigma}^{1/2}\left(\bm{\beta}_{0}\cdot\mathbb{I}_{\{t\leq\theta\}}+\bm{\beta}_{1}\cdot\mathbb{I}_{\{t>\theta\}}\right),$ (2)
@@
where $(\bm{\Sigma}^{1/2})^{\dagger}$ denotes the (Moore–Penrose) pseudoinverse of $\bm{\Sigma}^{1/2}$. Hence the lower bound reported in Table 1 is straightforward generalisation of Theorem 1 to the model class $\mathcal{P}^{\mathfrak{s},n,p}_{\bm{\Sigma},\sigma}(\tau)$ with general $\bm{\Sigma}$. It is readily seen that the phase transition is determined by the sparsity of the covariance weighted differential parameter, i.e. the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. Also, the lower bound is formulated with the size of the change measured by $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}$, complementing the following lemma:
Lemma 2 (Lemma 1 of Cho et al., 2025).
@@
Let $\mathsf{P}_{n}(\bm{\beta})$ denote the joint distribution of $\{(Y_{t},\mathbf{x}_{t}):\,t\in[n]\}$ such that $Y_{t}=\mathbf{x}_{t}^{\top}\bm{\beta}+\varepsilon_{t}$ for all $t$. We also denote by $\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})$ the joint distribution of $\{(Y_{t},\mathbf{x}_{t}):\,t\in[n]\}$ under (1). Then the total variation distance between $\mathsf{P}_{n}(\bm{\beta})$ and $\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})$, denoted by $\mathrm{TV}\bigl(\mathsf{P}_{n}(\bm{\beta}),\,\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})\bigr)$, satisfies
@@
$\frac{1}{100}\leq\frac{\min_{\bm{\beta},\bm{\beta}_{0},\bm{\beta}_{1}:\,\bm{\beta}_{1}-\bm{\beta}_{0}=\bm{\delta}}\mathrm{TV}\bigl(\mathsf{P}_{n}(\bm{\beta}),\,\mathsf{P}_{\theta}(\bm{\beta}_{0},\bm{\beta}_{1})\bigr)}{\min\left\{1,\,\sqrt{\frac{\theta(n-\theta)}{n\sigma^{2}}\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}}\right\}}\leq\frac{3\sqrt{3}}{2}.$
@@
Figure 2 numerically confirms these observations when $\bm{\Sigma}=[0.6^{|i-j|}]_{i,j}$. Specifically, the best performance of various change point estimation methods exhibits a steeper dependence on the intrinsic sparsity (middle panel) than on the standard sparsity (left panel), which is in alignment with the theoretical results reported in Table 1 (see also the right panel). Besides, we notice that McScan, a method attaining minimax rate-optimality in the sparse regime, has its estimation error exceed that of QcScan, a method tailored to achieve optimal performance in the dense regime, close to where the inherent sparsity $\mathfrak{s}$ meets $\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ (middle panel). On the other hand, this phenomenon occurs far before the sparsity of $\bm{\delta}$ reaches $\mathfrak{s}_{\delta}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$, which is accounted for by that $|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=p$ in all the scenarios considered in the left panel.
@@
Figure 2: Estimation performance of MOSEG (Cho and Owens, 2024), CHARCOAL (Gao and Wang, 2022), and the proposed methods: McScan (Section 3.1; a variant of Cho et al., 2025), QcScan (Section 3.2), OcScan (Section 4.1) and OcScan.R (Section 4.2), when $\bm{\Sigma}=[0.6^{|i-j|}]_{i,j}$. Taken from (M2) in Section 5, we set $n=300$, $p=200$, $\bm{\beta}_{0}+\bm{\beta}_{1}=\mathbf{0}$ and $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}=2$. In the left, the $x$-axis denotes the standard sparsity $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$ while in the middle, it is the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. In each scenario, results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The right panel plots the quantity $\min(|\bm{\Sigma}^{1/2}\bm{\delta}|_{0},\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}})$ as either $\mathfrak{s}_{\delta}$ or $\mathfrak{s}$ varies. This quantity is the influencing factor, save logarithmic factors, appearing in the minimax rate of estimation (see Table 1), and represents a proxy of the difficulty of the estimation problem. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ or $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@
In the high-dimensional change point detection literature, the phase transition with respect to sparsity has been noted in the context of mean change detection, see Enikeeva and Harchaoui (2019), Liu et al. (2021), Pilliat et al. (2023) and Kovács et al. (2024), all considering the case of cross-sectional independence. Theorem 1 is a first of its kind for the linear model setting where the cross-sectional dependence is present in $\mathbf{x}_{t}Y_{t}$ (whose expectation is piecewise constant as $\mathsf{E}(\mathbf{x}_{t}Y_{t})=\bm{\Sigma}\bm{\beta}_{0}+\bm{\Sigma}\bm{\delta}\cdot\mathbb{I}_{\{t\geq\theta+1\}}$), even if $\mathsf{Cov}(\mathbf{x}_{t})=\mathbf{I}$. Moreover, the sparsity inherent to this change point problem is characterised by $|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$ rather than the $\ell_{0}$-norm of $\bm{\delta}$, which also sets our findings apart from the existing work.
@@ 3.1 Sparse regime
Cho et al. (2025) propose McScan, a data segmentation method for the multiple change point problem in linear models. Motivated by the observation that $\mathsf{E}(\mathbf{x}_{t}Y_{t})$ contains the information about local parameter vector, it searches for change points by scanning the local covariance between $Y_{t}$ and $\mathbf{x}_{t}$. Avoiding the costly computation of local parameter estimates via $\ell_{1}$-regularision, it is highly efficient computationally and moreover, attains better statistical efficiency over a wider class of change point problems when compared to the state-of-the-art, by relaxing the conditions of exact sparsity imposed on the regression or differential parameters.
@@
We propose to evaluate $\bar{T}_{k}$ at strategically selected locations for the detection and estimation of the change point (if any). For this, we adopt the advanced optimistic search proposed by Kovács et al. (2024), see Algorithm 1 for the generic pseudo code applicable with any detector statistic denoted by $V_{k}$. It first evaluates $V_{k}$ on a dyadic grid denoted by $\mathcal{D}$, for a given trimming parameter $\varpi$, and identifies the maximiser $k^{*}$. If $V_{k^{*}}$ does not exceed some threshold $\zeta$, we quit the algorithm while if $V_{k^{*}}>\zeta$, we perform a divide-and-conquer-type search through calling the function OS over the interval determined with $k^{*}$ and its neighbours on the dyadic grid. Through this, we iteratively compare $V_{k}$ at dynamically selected search locations based on prior evaluations and discard the sections of the data that are not likely to contain any change point. Replacing the conventional exhaustive search, this adaptive strategy significantly reduces the total computational complexity from linear to logarithmic, while maintaining the statistical efficiency. We apply Algorithm 1 with McScan-specific detector statistic $\bar{T}_{k}$, threshold $\zeta_{\text{\tiny M}}$ and trimming parameter $\varpi_{\text{\tiny M}}$, to obtain the estimators of the number and location of the change point, $(\widehat{\theta}_{\text{\tiny M}},\widehat{q}_{\text{\tiny M}})$. With a slight abuse of terminology, we refer to the resulting method also as McScan (maximum covariance scanning).
@@
1ex
[…]
Algorithm 1 Advanced optimistic search (adapted from Kovács et al., 2024).
@@
$|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}\Delta\geq\bar{c}_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}\Psi^{2}\log(p\log(n)),$ (5)
or that there is no change point, with $q=0$ and $\theta=n$. Then, McScan applied with $\zeta_{\text{\tiny M}}=\bar{c}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\log(p\log(n))}$ and $\varpi_{\text{\tiny M}}$, outputs $(\widehat{\theta}_{\text{\tiny M}},\widehat{q}_{\text{\tiny M}})$ which satisfy
$\mathsf{P}\left\{\widehat{q}_{\text{\tiny M}}=q\,;\,|\widehat{\theta}_{\text{\tiny M}}-\theta|\leq\bar{c}_{1}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}\Psi^{2}\log(p\log(n))\right\}\geq 1-\bar{c}_{2}(p\log(n))^{-1}$
for some universal constants $\bar{c}$ and $\bar{c}_{i}\in(0,\infty),\,0\leq i\leq 2$.
@@
Thus, as long as $\sigma_{*}$ is bounded away from zero, McScan is minimax rate optimal up to a possible logarithmic factor in the sparse regime $\mathfrak{s}<\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ (cf. Table 1), without requiring any prior knowledge of $\bm{\Sigma}$, or ruling out strong collinearity among the variables that results in rank-deficient $\bm{\Sigma}$ — a situation that often arises with increasing number of variables. This is in stark contrast to the most existing methods, which require all eigenvalues of $\bm{\Sigma}$ to be uniformly bounded above and away from zero; see Section S1.1 of Cho et al. (2025) and also Figure 5 in Section 5 for numerical validation. When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ as in (2) allows for re-formulating the condition in (5), to
@@
Since $|\bm{\Sigma}^{1/2}\bm{\delta}|_{\infty}\geq\mathfrak{s}^{-1/2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}$ (recall that $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$), the above displayed condition is independent of $\sigma_{*}$, and matches the minimax lower bound on the detection boundary in the sparse regime provided that $\log(n)\lesssim p$ and $\mathfrak{s}$ is not very close to $\sqrt{p\log\log(n)}$.
@@
The original McScan algorithm of Cho et al. (2025) combines the detector statistic in (3) with the narrowest-over-threshold principle (Baranowski et al., 2019) and the seeded intervals (Kovács et al., 2023), evaluating the detector statistics over strategically selected intervals and selecting the significant local maximisers for the detection of multiple change points. Theorem 3 is a variant of the more general Theorem 2 in Cho et al. (2025) which, permitting serial dependence and sub-Weibull tail behaviour in the data, shows the consistency of McScan in multiple change point detection, all without placing any condition on the exact sparsity of regression or differential parameters, or that on $p$ beyond what is implied by the detection lower bound comparable to (5). There, it is further demonstrated that McScan allows for a strictly wider class of change point models in comparison with the state-of-the-art.